OpenStudy (pre-calculus):

can someone walk me through the steps of how to simplify 2x(3x)^2

OpenStudy (joshtray):

You mean\[2x(3x)^{2}\] right?

OpenStudy (welshfella):

deal with the exponent part first (3x)^2 = 3x * 3x = 3^2 x^2 = 9x^2

OpenStudy (pre-calculus):

@Joshtray yes

OpenStudy (welshfella):

now multiply 9x^2 by 2x

OpenStudy (joshtray):

\[9x ^{2}\times2x=18x ^{3}\]

OpenStudy (joshtray):

So therefore our final answer is 18x^3

OpenStudy (pre-calculus):

alright thank you so much guys!! @Joshtray @welshfella

OpenStudy (joshtray):

You're welcome

OpenStudy (welshfella):

you need to remember the rules of exponents 1. x^a * x^b = x^(a+b) 2. x^a / x^b = x ^(a-b) 3. (x^a)^b = x^(ab)

OpenStudy (welshfella):

That is Multiplying - you add exponents Dividing you subtract exponents Taking one exponent to another power you multiply exponents.